OpenStudy (ashy98):

Find the length of a rectangle that has a width of 0.8 cm and area of 3.2 cm^2

OpenStudy (anonymous):

Area = Length times the width so make an Equation A=L*W A=3.2 W=0.8

OpenStudy (anonymous):

3.2=L*0.8

OpenStudy (anonymous):

Divide both sides by 0.8

OpenStudy (anonymous):

If you do that, you get L=4

OpenStudy (anonymous):

The answer is 4 cm

OpenStudy (ashy98):

thank you, you help me alot to know what im doing lol (: